# Hybrid estimation for a mixed fractional Black-Scholes model with random effects from discrete time observations

### v3

@@ 7.8 Proof of Proposition 6
we obtain $\sum\limits_{\underset{j<l}{j,l=1}}^{m}F(x_{j})\mathcal{L}_{j}\left(x\right)\mathcal{L}_{l}\left(x\right)=\dfrac{1}{2}F(x)\left(1-S_{m}(x)\right)+O(m^{-1}).$ Finally, by replacing this term in (7.16), we conclude for $x\in[-1,1]$,
$\lim\limits_{n,N\to\infty}\mathrm{Var}\left(\hat{F}_{m,n,N}(x)\right)=\lim\limits_{n,N\to\infty}\dfrac{1}{N}\mathbb{E}\left(Z_{1,N}^{2}\right)=\dfrac{1}{N}F(x)\left(1-F(x)\right)+O(N^{-1}m^{-1/2}),$
The third assertion of Proposition 6 is a deduction from the two previous assertion. Indeed, for $x\in[-1,1]$, the MSE of $\hat{F}_{m,n,N}$ is defined as follows
$\mathrm{MSE}\left(\hat{F}_{m,n,N}(x)\right)$ $:=$ $\displaystyle\mathbb{E}\left[\left(\hat{F}_{m,n,N}(x)-F(x)\right)^{2}\right]$
$=$ $\mathrm{Var}\left(\hat{F}_{m,n,N}(x)\right)+\left(Bias\left(\hat{F}_{m,n,N}(x)\right)\right)^{2},$
and its asymptotic value, which we denote by $\mathrm{AMSE}\left(\hat{F}_{m,n,N}(x)\right)$, can be written as the following
$\mathrm{AMSE}\left(\hat{F}_{m,n,N}(x)\right)=\underset{n,N\to\infty}{\lim}\mathrm{Var}\left(\hat{F}_{m,n,N}(x)\right)+\underset{n,N\to\infty}{\lim}\left(Bias\left(\hat{F}_{m,n,N}(x)\right)\right)^{2}.$
@@
$\mathrm{AMSE}\left(\hat{F}_{m,n,N}(x)\right)=\dfrac{1}{N}\sigma^{2}_{F}(x)+\frac{\pi^{2}}{m^{4}}\left(T_{m}(x)\mu(x)\right)^{2}+O\left(N^{-1}m^{-1/2}\right)+O(m^{-4}),$
@@ 7.9 Proof of Theorem 2
For $x\in[-1,1]$, the bias-variance decomposition of the error of $\hat{F}_{m,n,N}$ combined with the triangle inequality allows us to write
$\left\|\hat{F}_{m,n,N}-F\right\|\leq\left\|\hat{F}_{m,n,N}-\mathbb{E}\left(\hat{F}_{m,n,N}\right)\right\|+\left\|\mathbb{E}\left(\hat{F}_{m,n,N}\right)-F\right\|.$ (7.19)
Let us begin by analyzing the bias term $\left\|\mathbb{E}\left(\hat{F}_{m,n,N}\right)-F\right\|$. For $x\in(-1,1]$, we define
@@
$\tilde{F}_{m}(x):=\mathbb{E}\left(\hat{F}_{m,n,N}\right)=\sum\limits_{j=1}^{m}\mathbb{P}\left(\widehat{\phi}_{i}\leq x_{j}\right)\mathcal{L}_{j}(x)\ \text{and}\ F_{m}(x):=\sum\limits_{j=1}^{m}F(x_{j})\mathcal{L}_{j}(x).$
Using the triangle inequality again, we decompose the bias term as follows
$\left\|\tilde{F}_{m}-F\right\|\leq\left\|\tilde{F}_{m}-F_{m}\right\|+\left\|F_{m}-F\right\|.$ (7.20)
For the first term of the right-hand side of (7.20), we have for $x\in[-1,1]$,
$\left|\tilde{F}_{m}(x)-F_{m}(x)\right|$ $\leq$ $\sum\limits_{j=1}^{m}\left|F(x_{j})-\mathbb{P}\left(\hat{\phi}_{1}\leq x_{j}\right)\right|\left|\mathcal{L}(x_{j})\right|$
$\leq$ $\underset{1\leq j\leq m}{\max}\left|F(x_{j})-\mathbb{P}\left(\hat{\phi}_{1}\leq x_{j}\right)\right|\ \sum\limits_{j=1}^{m}\left|\mathcal{L}(x_{j})\right|.$
Since for each $i=1,\ldots,N$, $\hat{\phi}_{i}\overset{a.s}{\longrightarrow}\phi_{i}$, as $n,N\to\infty$, the Portmanteau theorem ensures that, for each fixed $x_{j}$,
$\mathbb{P}\left(\hat{\phi}_{i}\leq x_{j}\right)\to F(x_{j}),\quad\text{for each}\ i=1,\ldots,N,$
which implies that
$\left|F(x_{j})-\mathbb{P}\left(\hat{\phi}_{1}\leq x_{j}\right)\right|\longrightarrow 0,\quad\text{for each}\ j=1,\ldots,m.$
To handle the supremum over $x$, we use the Lebesgue constant property. Since $\sum_{j=1}^{m}|L_{j}(x)|\leq\Lambda_{m}$ uniformly in $x\in[-1,1]$, where $\Lambda_{m}=O(\log m)$ is the Lebesgue constant for Chebyshev nodes, we obtain
$\|\tilde{F}_{m}-F_{m}\|$ $=\sup_{x\in[-1,1]}|\tilde{F}_{m}(x)-F_{m}(x)|$
$\leq\sup_{x\in[-1,1]}\max_{1\leq j\leq m}|F(x_{j})-P(\hat{\phi}_{1}\leq x_{j})|\sum_{j=1}^{m}|L_{j}(x)|$
$\leq\max_{1\leq j\leq m}|F(x_{j})-P(\hat{\phi}_{1}\leq x_{j})|\sup_{x\in[-1,1]}\sum_{j=1}^{m}|L_{j}(x)|$
$\leq\Lambda_{m}\max_{1\leq j\leq m}|F(x_{j})-P(\hat{\phi}_{1}\leq x_{j})|.$
Since $\Lambda_{m}=O(\log m)$ is bounded for fixed $m$, and each term $|F(x_{j})-P(\hat{\phi}_{1}\leq x_{j})|\to 0$ as $n,N\to\infty$, we conclude that for fixed $m$, we obtain
$\|\tilde{F}_{m}-F_{m}\|\to 0,\quad\mathrm{as}\quad n,N\to\infty.$ (7.21)
For the second term of the right hand side of (7.20), since $F\in\mathcal{C}^{2}\left([-1,1]\right)$, the Jackson’s theorem implies that
$\underset{x\in[-1,1]}{\sup}\left|F_{m}(x)-F(x)\right|\ \leq\dfrac{C}{m^{2}},\quad\text{for some constant}\ C>0.$
Then, when $m$ tends to infinity, we obtain
$\left\|F_{m}-F\right\|\to 0.$ (7.22)
Finally, combining (7.21) with (7.22), we conclude that
$\left\|\tilde{F}_{m}-F\right\|\ \xrightarrow{a.s}0,\quad\text{as}\ n,N\to\infty\ \text{and then}\ m\to\infty.$ (7.23)
We now proceed to analyze the variance term $\|\hat{F}_{m,n,N}-\tilde{F}_{m}\|$. Since $\hat{F}_{m,n,N}(x)$ is the Lagrange interpolation of $\hat{F}_{N}$ at the Chebyshev nodes, and using properties of Lagrange interpolation on [1,1][-1,1], for each $x\in[-1,1]$, we have
$\hat{F}_{m,n,N}(x)-\tilde{F}_{m}(x)$ $=\sum_{j=1}^{m}[\hat{F}_{N}(x_{j})-E[\hat{F}_{N}(x_{j})]]L_{j}(x)$ (7.24)
$=\sum_{j=1}^{m}[\hat{F}_{N}(x_{j})-P(\hat{\phi}_{i}\leq x_{j})]L_{j}(x).$ (7.25)
Taking the supremum over $x\in[-1,1]$ and using the Lebesgue constant, we get
$\|\hat{F}_{m,n,N}-\tilde{F}_{m}\|$ $\leq\sup_{x\in[-1,1]}\sum_{j=1}^{m}|\hat{F}_{N}(x_{j})-P(\hat{\phi}_{i}\leq x_{j})||L_{j}(x)|$ (7.26)
$\leq\max_{1\leq j\leq m}|\hat{F}_{N}(x_{j})-P(\hat{\phi}_{i}\leq x_{j})|\sup_{x\in[-1,1]}\sum_{j=1}^{m}|L_{j}(x)|$ (7.27)
$\leq\Lambda_{m}\max_{1\leq j\leq m}|\hat{F}_{N}(x_{j})-P(\hat{\phi}_{i}\leq x_{j})|.$ (7.28)
By using the Glivenko-Cantelli theorem applied to the finite set of points $\{x_{1},\ldots,x_{m}\}$, we obtain
$\max_{1\leq j\leq m}|\hat{F}_{N}(x_{j})-P(\hat{\phi}_{i}\leq x_{j})|\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0,\text{ as }N\to\infty.$
Since $\Lambda_{m}=O(\log m)$ grows slowly, we have
$\|\hat{F}_{m,n,N}-\tilde{F}_{m}\|\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0,\text{ as }N\to\infty.\quad$ (7.29)
Finally, combining (7.23) with (7.29) under the condition $m=o\big(\frac{\sqrt{N}}{\log N}\big)$, we obtain the uniform convergence of $\hat{F}_{m,n,N}$.
@@ 7.10 Proof of Theorem 3
For $x\in[-1,1]$, we split the error term as follows
$\sqrt{N}\left(\hat{F}_{m,n,N}(x)-F(x)\right)=\sqrt{N}\left(\hat{F}_{m,n,N}(x)-\hat{F}_{N}(x)\right)+\sqrt{N}\left(\hat{F}_{N}(x)-F(x)\right),$ (7.30)
where $\hat{F}_{N}(x)=\dfrac{1}{N}\sum\limits_{i=1}^{N}\mathds{1}_{\{\hat{\phi}_{i}\leq x\}}$. The first term of (7.30), the estimator $\hat{F}_{m,n,N}$ can be seen as the Lagrange interpolation of $\hat{F}_{N}$ on [1,1][-1,1]. This allows us to write for $x\in[-1,1]$ and some positive constant $C$,
$\left|\hat{F}_{m,n,N}(x)-\hat{F}_{N}(x)\right|\leq Cm^{-2}.$ (7.31)
Scaling both sides by $\sqrt{N}$, we obtain for $x\in[-1,1]$,
$\sqrt{N}\left|\hat{F}_{m,n,N}(x)-\hat{F}_{N}(x)\right|\leq CN^{\frac{1}{2}}m^{-2}.$
If we choose $m$ such that $N^{\frac{1}{2}}m^{-2}\to 0$ as $N\to\infty$, we deduce that
$\sqrt{N}\left(\hat{F}_{m,n,N}(x)-\hat{F}_{N}(x)\right)\xrightarrow{\mathbb{P}}0,\quad\text{as}\ N\to\infty.$ (7.32)
We now analyze the second term of (7.30). For each $x\in[-1,1]$, we write
$\sqrt{N}\left(\hat{F}_{N}(x)-F(x)\right)=\sqrt{N}\left(\hat{F}_{N}(x)-\mathbb{P}(\hat{\phi}_{i}\leq x)\right)+\sqrt{N}\left(\mathbb{P}(\hat{\phi}_{i}\leq x)-F(x)\right).$
Since $\hat{F}_{N}(x)$ is the empirical distribution function of estimators $\hat{\phi}_{i}$, $i=1,\ldots,N$ which are i.i.d., it follows by the CLT that for $x\in[-1,1]$,
$\sqrt{N}\left(\hat{F}_{N}(x)-\mathbb{P}(\hat{\phi}_{i}\leq x)\right)\xrightarrow{d}\mathcal{N}\left(0,\ \mathbb{P}\left(\hat{\phi}_{i}\leq x\right)\left(1-\mathbb{P}\left(\hat{\phi}_{i}\leq x\right)\right)\right),\ \ \text{as}\ \ N\to\infty.$ (7.33)
Moreover, since for each $i=1,\ldots,N$, $\hat{\phi}_{i}$ is a strongly consistent estimator of $\phi_{i}$ when $n,N$ tend to infinity, it follows by the Portmanteau theorem that for $x\in[-1,1]$,
$\mathbb{P}\left(\hat{\phi}_{i}\leq x\right)\xrightarrow{a.s}F(x),\quad\text{as}\quad n,N\to\infty,$
which implies that for $x\in[-1,1]$,
$\sqrt{N}\left(\mathbb{P}(\hat{\phi}_{i}\leq x)-F(x)\right)\xrightarrow{\mathbb{P}}0,\quad n,N\to\infty.$ (7.34)
Combining (7.33) with (7.34), the Slustky’s theorem ensures that for $x\in[-1,1]$
$\sqrt{N}\left(\hat{F}_{N}(x)-F(x)\right)\xrightarrow{d}\mathcal{N}\left(0,F(x)\left(1-F(x)\right)\right).$ (7.35)
Finally, combining (7.32) with (7.35), we conclude by Slutsky’s theorem,
$\sqrt{N}\left(\hat{F}_{m,n,N}(x)-F(x)\right)\xrightarrow{d}\mathcal{N}\left(0,F(x)\left(1-F(x)\right)\right),$
which proves the asymptotic normality of $\hat{F}_{m,n,N}$.

### v4

@@ 7.8 Proof of Proposition 6
we obtain $\sum\limits_{\underset{j<l}{j,l=1}}^{m}F(x_{j})\mathcal{L}_{j}\left(x\right)\mathcal{L}_{l}\left(x\right)=\dfrac{1}{2}F(x)\left(1-S_{m}(x)\right)+O(m^{-1}).$
Finally, by replacing this term in (7.25), we conclude for $x\in[-1,1]$,
$\lim_{\begin{subarray}{c}n,N\to\infty\end{subarray}}\dfrac{1}{N}\mathbb{E}\left(Z_{1,N}^{2}\right)=\dfrac{1}{N}F(x)\left(1-F(x)\right)+O(N^{-1}m^{-\frac{1}{2}}).$
It remains to control the covariance term of (7.24). Using Lemma (3), we obtain
$\frac{2}{N^{2}}\sum_{1\leq i<k\leq N}\mathrm{Cov}(Z_{i,N},Z_{k,N})$ $\leq$ $\frac{2}{N^{2}}\sum_{1\leq i<k\leq N}\left|\mathrm{Cov}(Z_{i,N},Z_{k,N})\right|$
$\leq$ $\frac{2}{N^{2}}\sum_{1\leq i<k\leq N}\frac{C}{N}=\frac{N\big(N-1\big)}{2}\,\frac{C}{N},$
which implies that $\dfrac{2}{N^{2}}\sum\limits_{1\leq i<k\leq N}\mathrm{Cov}(Z_{i,N},Z_{k,N})=O(N^{-1})$.
Combining the above results, we conclude that
$\mathrm{Var}\left(\widehat{F}_{m,n,N}(x)\right)=\frac{1}{N}F(x)(1-F(x))+O(N^{-1}m^{-1/2}),$
which proves the second assertion of Proposition 6. The third assertion of Proposition 6 is a deduction from the two previous assertion. Indeed, for $x\in[-1,1]$, the MSE of $\widehat{F}_{m,n,N}$ is defined as follows
$\mathrm{MSE}\left(\widehat{F}_{m,n,N}(x)\right)$ $:=$ $\displaystyle\mathbb{E}\left[\left(\widehat{F}_{m,n,N}(x)-F(x)\right)^{2}\right]$
$=$ $\mathrm{Var}\left(\widehat{F}_{m,n,N}(x)\right)+\left(Bias\left(\widehat{F}_{m,n,N}(x)\right)\right)^{2},$
and its asymptotic value, which we denote by $\mathrm{AMSE}\left(\widehat{F}_{m,n,N}(x)\right)$, can be written as the following
$\mathrm{AMSE}\left(\widehat{F}_{m,n,N}(x)\right)=\underset{n,N\to\infty}{\lim}\mathrm{Var}\left(\widehat{F}_{m,n,N}(x)\right)+\underset{n,N\to\infty}{\lim}\left(Bias\left(\widehat{F}_{m,n,N}(x)\right)\right)^{2}.$
@@
$\mathrm{AMSE}\left(\widehat{F}_{m,n,N}(x)\right)=\dfrac{1}{N}\sigma^{2}_{F}(x)+\frac{\pi^{2}}{m^{4}}\left(T_{m}(x)\mu(x)\right)^{2}+O\left(N^{-1}m^{-\frac{1}{2}}\right)+O(m^{-4}),$
@@ 7.9 Proof of Theorem 2
To prove the consistency of $\widehat{F}_{m,n,N}$, we use the following decomposition
$\|\widehat{F}_{m,n,N}-F\|\leq\underbrace{\|\widehat{F}_{m,n,N}-\widehat{F}_{N}\|}_{\text{(A)}}+\underbrace{\|\widehat{F}_{N}-\bar{F}_{N}\|}_{\text{(B)}}+\underbrace{\|\bar{F}_{N}-F\|}_{\text{(C)}},$ (7.28)
where
$\widehat{F}_{N}(x):=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{\widehat{\phi}_{i}\leq x\}},\qquad\bar{F}_{N}(x):=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{\phi_{i}\leq x\}}.$
We shall show that each term converges to zero almost surely along a suitable diagonal $n=n_{N}$.
We begin with the term $(C)$. Since the random effects $\phi_{i}$, $i=1,\ldots,N$ are i.i.d, the classical Glivenko-Cantelli theorem ensures that
$\|\bar{F}_{N}-F\|\overset{a.s}{\longrightarrow}0,\qquad N\to\infty.$ (7.29)
Let us now control the term $(B)$. Fix $\varepsilon>0$. For any $x\in[-1,1]$ and any $i\leq N$, we have
$|\mathds{1}_{\{\widehat{\phi}_{i}\leq x\}}-\mathds{1}_{\{\phi_{i}\leq x\}}|\leq\mathds{1}_{\{|\widehat{\phi}_{i}-\phi_{i}|>\varepsilon\}}+\mathds{1}_{\{x-\varepsilon<\phi_{i}\leq x+\varepsilon\}}.$
Averaging over $i$ and taking the supremum over $x\in[-1,1]$ yields
$\|\widehat{F}_{N}-\bar{F}_{N}\|\leq\,\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{|\widehat{\phi}_{i}-\phi_{i}|>\varepsilon\}}+\sup_{x\in[-1,1]}\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{x-\varepsilon<\phi_{i}\leq x+\varepsilon\}},$ (7.30)
Since $\widehat{\phi}_{i}=\widehat{\theta}_{i}+\frac{1}{2}\widehat{\sigma}^{2}$ and $\phi_{i}=\theta_{i}+\frac{1}{2}\sigma^{2}$, we have
$|\widehat{\phi}_{i}-\phi_{i}|\leq|\widehat{\theta}_{i}-\theta_{i}|+\frac{1}{2}|\widehat{\sigma}^{2}-\sigma^{2}|.$
For each fixed $i\leq N$, Proposition 1 gives $\widehat{\theta}_{i}\xrightarrow{a.s.}\theta_{i}$ as $n\to\infty$. Moreover, from Proposition 4, $\widehat{\sigma}^{2}\xrightarrow{a.s.}\sigma^{2}$ as $n\to\infty$ first, then $N\to\infty$. Let $\varepsilon>0$. Then, for each fixed $N$, there exists $n_{0}=n_{0}(N,\varepsilon)$ such that for all $n\geq n_{0}$,
@@
$\mathbb{P}\left(\max\limits_{1\leq i\leq N}|\widehat{\theta}_{i}-\theta_{i}|>\frac{\varepsilon}{2}\right)\leq\frac{1}{4N^{2}},$
and
$\mathbb{P}\left(|\widehat{\sigma}^{2}-\sigma^{2}|>\varepsilon\right)\leq\frac{1}{4N^{2}}.$ (7.31)
Define the sequence $n=n_{N}:=n_{0}(N,\varepsilon_{N})$ where $\varepsilon_{N}=N^{-\frac{1}{2}}$ (for example). Then by the union bound,
$\mathbb{P}\left(\max\limits_{1\leq i\leq N}|\widehat{\phi}_{i}-\phi_{i}|>\varepsilon_{N}\right)$ $\leq\mathbb{P}\left(\max\limits_{1\leq i\leq N}|\widehat{\theta}_{i}-\theta_{i}|>\frac{\varepsilon_{N}}{2}\right)+\mathbb{P}\left(|\widehat{\sigma}^{2}-\sigma^{2}|>\varepsilon_{N}\right)$
$\leq\frac{1}{4N^{2}}+\frac{1}{4N^{2}}=\frac{1}{2N^{2}}.$
Since $\sum\limits_{N=1}^{\infty}\frac{1}{2N^{2}}<\infty$, the Borel-Cantelli lemma implies that a.s, the event $\left\{\max\limits_{1\leq i\leq N}|\widehat{\phi}_{i}-\phi_{i}|>\varepsilon_{N}\right\}$ occurs only finitely many times. In other words, there exists a integer $N_{0}<\infty$ such that for all $N\geq N_{0}$,
$\max\limits_{1\leq i\leq N}|\widehat{\phi}_{i}-\phi_{i}|\leq\varepsilon_{N},\quad\text{a.s.}$
Therefore, as $N\to\infty$, we obtain
$\max\limits_{1\leq i\leq N}|\widehat{\phi}_{i}-\phi_{i}|\leq\varepsilon_{N}\overset{a.s}{\longrightarrow}0.$ (7.32)
Hence, for every fixed $\varepsilon>0$,
$\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{|\widehat{\phi}_{i}-\phi_{i}|>\varepsilon\}}=0\quad\text{a.s}\quad\text{for all sufficiently large }\,N.$ (7.33)
For the second term of the right-hand side of (7.30), by the strong law of large numbers applied to the i.i.d. sample $\phi_{i}$, $i=1,\ldots,N$, we obtain for each fixed $x\in[-1,1]$,
$\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{x-\varepsilon<\phi_{i}\leq x+\varepsilon\}}\xrightarrow{a.s.}\mathbb{P}(x-\varepsilon<\phi\leq x+\varepsilon).$
Moreover, by assumption (A3), we get
$\mathbb{P}(x-\varepsilon<\phi\leq x+\varepsilon)\leq\int_{x-\varepsilon}^{x+\varepsilon}f(u)\,du\leq 2\|f\|\,\varepsilon.$
Consequently,
$\limsup_{N\to\infty}\sup_{x\in[-1,1]}\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{x-\varepsilon<\phi_{i}\leq x+\varepsilon\}}\leq 2\|f\|\,\varepsilon,\qquad\text{a.s.}$ (7.34)
Replacing (7.33) and (7.34) into (7.30) gives
$\limsup_{N\to\infty}\|\widehat{F}_{N}-\bar{F}_{N}\|\leq 2\|f\|\,\varepsilon\qquad\text{a.s.}$ (7.35)
Since $\varepsilon>0$ was arbitrary, then letting $\varepsilon\rightarrow 0$ yields
$\|\widehat{F}_{N}-\bar{F}_{N}\|\xrightarrow{a.s.}0\qquad\text{as }N\to\infty\text{ along }n=n_{N}.$ (7.36)
We now proceed to control the first term of (7.28). By Lebesgue’s Lemma, we have
$\|\widehat{F}_{m,n,N}-\widehat{F}_{N}\|\leq(1+\Lambda_{m})E_{m}(\widehat{F}_{N}),$
where $\Lambda_{m}=\sup\limits_{x\in[-1,1]}\sum\limits_{j=1}^{m}|L_{j}(x)|$ is the Lebesgue constant and $E_{m}(\widehat{F}_{N})=\inf\limits_{p\in\mathbb{P}_{m-1}}\|\widehat{F}_{N}-p\|$ is the best uniform approximation error by polynomials of degree at most m1m-1. For Chebyshev nodes, it is well‑known that $\Lambda_{m}=O(\log m)$. Since $F\in\mathcal{C}^{2}([-1,1])$, Jackson’s inequality gives
$E_{m}(F)\leq\frac{C_{1}}{m^{2}}\|f^{\prime}\|.$
From (7.36) and (7.29) we already know that $\|\widehat{F}_{N}-F\|\to 0$ along $n=n_{N}$. Hence, for large $N$,
$E_{m}(\widehat{F}_{N})\leq E_{m}(F)+\|\widehat{F}_{N}-F\|\leq\frac{C_{2}}{m^{2}}.$
Therefore
$\|\widehat{F}_{m,n,N}-\widehat{F}_{N}\|\leq\Big(1+C_{2}\log(m)\Big)\frac{C_{1}\|f^{\prime}\|}{m^{2}}\leq C^{\prime}\frac{\log m}{m^{2}}.$ (7.37)
If we choose $m=m_{N}$ satisfying $\frac{\log(m_{N})}{m_{N}^{2}}\to 0$ as $N\to\infty$, we obtain
$\|\widehat{F}_{m,n,N}-\widehat{F}_{N}\|\overset{a.s}{\longrightarrow}0\qquad\text{as }\quad N\to\infty.$ (7.38)
Replacing (7.29), (7.36) and (7.38) into the decomposition (7.28) yields
$\|\widehat{F}_{m_{N},n_{N},N}-F\|\to 0,\qquad N\to\infty.$
This completes the proof.
@@ 7.10 Proof of Theorem 3
For each $x\in[-1,1]$, we decompose the error term as follows
$\sqrt{N}\bigl(\widehat{F}_{m,n,N}(x)-F(x)\bigr)=\sqrt{N}\bigl(\widehat{F}_{m,n,N}(x)-\widehat{F}_{N}(x)\bigr)+\sqrt{N}\bigl(\widehat{F}_{N}(x)-F(x)\bigr).$ (7.39)
From equation (7.37) in the proof of Theorem 2, the Lagrange interpolation error satisfies
$\|\widehat{F}_{m,n,N}-\widehat{F}_{N}\|\;\leq\;C^{\prime}\frac{\log m}{m^{2}}.$
Hence,
$\sqrt{N}\left|\widehat{F}_{m,n,N}(x)-\widehat{F}_{N}(x)\right|\leq\sqrt{N}\,\|\widehat{F}_{m,n,N}-\widehat{F}_{N}\|\leq C^{\prime}\sqrt{N}\,\frac{\log m}{m^{2}}\;\xrightarrow{}0,$
Taking $m=m_{N}$ that verifies $\sqrt{N}\,\frac{\log m}{m^{2}}\;\xrightarrow{}0$, as $N\to\infty$, we obtain,
$\sqrt{N}\bigl(\widehat{F}_{m,n,N}(x)-\widehat{F}_{N}(x)\bigr)\xrightarrow{\mathbb{P}}0.$ (7.40)
Furthermore, since the variables $\widehat{\phi}_{i}$ are not independent, we write
$\widehat{\phi}_{i}=\phi_{i}+\big(\widehat{\theta}_{i}-\theta_{i}\big)+\tfrac{1}{2}(\widehat{\sigma}^{2}-\sigma^{2}).$
From Proposition 1 and Theorem 1, we have under the sequential regime,
$\max\limits_{1\leq i\leq N}|\widehat{\theta}_{i}-\theta_{i}|\xrightarrow{a.s.}0,\qquad\delta:=\tfrac{1}{2}(\widehat{\sigma}^{2}-\sigma^{2})=O_{\mathbb{P}}(N^{-1/2}).$
Using this decomposition, we may write
$\widehat{F}_{N}(x)=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{\phi_{i}\leq x-\delta-(\widehat{\theta}_{i}-\theta_{i})\}}=\bar{F}_{N}(x-\delta)+r_{N}(x),$
where $\bar{F}_{N}$ denotes the empirical distribution function of $\phi_{i}$, and where the remainder $r_{N}(x)$ satisfies
$\sqrt{N}\,r_{N}(x)\xrightarrow{P}0,$
uniformly in $x\in[-1,1]$. Hence,
$\widehat{F}_{N}(x)=\bar{F}_{N}(x-\delta)+o_{\mathbb{P}}(N^{-1/2}).$ (7.41)
Conditionally on $\delta$, the variables $\mathds{1}_{\{\phi_{i}\leq x-\delta\}}$ are i.i.d. Bernoulli with parameter $F(x-\delta)$. Hence, by the classical CLT,
$\sqrt{N}\bigl(\bar{F}_{N}(x-\delta)-F(x-\delta)\bigr)\xrightarrow{d}\mathcal{N}\!\left(0,F(x-\delta)\bigl(1-F(x-\delta)\bigr)\right).$ (7.42)
Since $F\in C^{2}([-1,1])$ and $\delta\to 0$ in probability, we have
$F(x-\delta)\to F(x),\qquad F(x-\delta)\bigl(1-F(x-\delta)\bigr)\to F(x)\bigl(1-F(x)\bigr),$ (7.43)
in probability. Moreover, a Taylor–Young expansion of $F$ around $x$ yields
$F(x-\delta)-F(x)=-f(x)\delta+o(|\delta|),$
so that, multiplying by $\sqrt{N}$ and using $\delta=O_{\mathbb{P}}(N^{-\frac{1}{2}})$,
$\sqrt{N}\bigl(F(x-\delta)-F(x)\bigr)=O_{\mathbb{P}}(1).$ (7.44)
Therefore,
$\sqrt{N}\bigl(\bar{F}_{N}(x-\delta)-F(x)\bigr)=\sqrt{N}\bigl(\bar{F}_{N}(x-\delta)-F(x-\delta)\bigr)+\sqrt{N}\bigl(F(x-\delta)-F(x)\bigr).$
By (7.42), the first term converges in distribution to a centered normal random variable with variance $F(x)\bigl(1-F(x)\bigr)$ (using (7.43)). The second term is $O_{\mathbb{P}}(1)$ by (7.44). Hence, Slutsky’s theorem implies that
$\sqrt{N}\bigl(\bar{F}_{N}(x-\delta)-F(x)\bigr)\xrightarrow{d}\mathcal{N}\!\left(0,F(x)\bigl(1-F(x)\bigr)\right).$ (7.45)
Using (7.41), we conclude that
$\sqrt{N}\bigl(\widehat{F}_{N}(x)-F(x)\bigr)\xrightarrow{d}\mathcal{N}\!\left(0,F(x)(1-F(x))\right).$ (7.46)
Finally, combining (7.39), (7.40), and (7.46) and applying Slutsky’s theorem once more, we obtain
$\sqrt{N}\bigl(\widehat{F}_{m,n,N}(x)-F(x)\bigr)\xrightarrow{d}\mathcal{N}\!\left(0,F(x)(1-F(x))\right),$
which completes the proof.